# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 4.2. Equivariant surgery preserving the fixed set
We make a rising waters argument; compare with [GS00], Prop. 6.2.1 for this type of argument for surfaces in 4-manifolds. Consider the equivariant radial Morse function $h$ on $B^{5}$ from Lemma 4.3. Note that both $Y$ and $\operatorname{Fix}(\tau)$ are properly embedded and boundary parallel, and their intersection is a single properly embedded tangle. Furthermore, $\operatorname{Fix}(\tau)$ consists of exactly one index 0 critical point with respect to $h$, and so does the intersection $\cap\operatorname{Fix}(\tau). YY$ itself has a unique index 0 critical point, but may have other critical points as well.
The minima of $Y$ and $\operatorname{Fix}(\tau)$ with respect to this Morse function both contribute a 1-handle each to the complement. The minimum point of their intersection then contributes a 2-handle, the attaching sphere of which winds around the commutator of the meridians of both $Y$ and $\operatorname{Fix}(\tau)$ near the intersection point. Thus the fundamental group of the sublevel set of $h$ just above the first intersection point is $\mathbb{Z}^{2}$. However, all subsequent critical points of $Y\cup\operatorname{Fix}(\tau)$ are in $Y$ and are of index 1 or 2, which contribute 2- or 3-handles to the complement. Of these, only the 2-handles can contribute to the fundamental group relations of $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, but these only introduce extra commutation relations between meridians that already commute. ∎
We now pass to the quotient under $\tau$, in which the meridians of $Y$ and $\operatorname{Fix}(\tau)$ are mapped to meridians of $\overline{Y}$ and $\operatorname{Branch}(\tau)$, respectively, and no new relations are added. Thus any curve that is null-homologous in $\overline{B^{5}}\setminus(Y\cup\operatorname{Branch}(\tau))$ is also null-homotopic in $\overline{B^{5}}\setminus(\overline{Y}\cup\operatorname{Branch}(\tau))$. Hence we conclude that if $\gamma$ is a curve in $\overline{Y}$ that is null-homologous in $\overline{B^{5}}\setminus\operatorname{Branch}(\tau)$, then it bounds a disk in $\overline{B^{5}}\setminus(\overline{Y}\cup\operatorname{Branch}(\tau))$ and we may therefore surger along it.
4.3. Finding surgery curves
Now to ambiently surger $(\overline{Y},\overline{A})$ to $(\overline{B^{3}},\overline{\ell})$, we proceed in two steps. The first is to ignore $\overline{A}$ and simply find surgery curves that realize a spin cobordism between $\overline{Y}$ and $\overline{B^{3}}$. We can find such curves by Kaplan’s theorem [Kap79]. Once we identify the curves we wish to surger $\overline{Y}$ along, we isotope them inside $\overline{Y}$ so that they are null-homologous in $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, allowing them to pass through $\overline{A}$ if necessary. Such an isotopy inside $\overline{Y}$ can always be found as $\overline{A}$ is null-homologous in $\overline{Y}$ rel. boundary, and so if a surgery curve $\gamma$ in $\overline{Y}$ does not bound a disk in $\overline{B^{5}}\setminus\overline{Y}\cup\operatorname{Fix}(\tau)$, we may use a surface in $\overline{Y}$ bounded by $\overline{A}$ to isotope it along a path to $\overline{A}$ and push it past $\overline{A}$ to remove the intersection. The result of the surgery on these curves will be to turn $\overline{Y}$ into $\overline{B^{3}}$, realizing $\overline{A}$ as an arc in a 3-ball.
It remains then to unknot $\overline{A}$ via surgery preserving $B^{3}$. We do this diagrammatically with explicit surgery curves that are null-homologous in the complement of $\overline{A}$. To do this, choose an arc in $S$ connecting the two boundary points of $\overline{A}$, and let $B^{\prime}$ be the knot obtained as the union of this arc with $\overline{A}$, pushed into the interior of $B^{3}$. Recall by Lemma 3.5 that if $\operatorname{Arf}(B^{\prime})=0$, then $B^{\prime}$ is band pass equivalent to the unknot, and that these band passes can be achieved by 0-framed (thus spin) surgery on null-homologous curves, as shown in Figure 9. If $\operatorname{Arf}(B^{\prime})=1$, then we can use Lemma 3.2 to surger $\overline{A}$ to the trefoil tangle if we wish.
@@ 4.3. Finding surgery curves
[…]
Figure 9. A band pass achieved by 0-framed 2-handles.
4.4. Lifting the surgeries
Let $\{\gamma_{i},\Delta_{i}\}$ be the collection of framed curves $\gamma$ in $\overline{Y}$ that we wish to surger along, and let $\Delta_{i}$ be the framed disks that they bound, as found in the previous subsection. Then since all the $\gamma_{i}$ are null-homotopic in the complement of the branch set, they lift to $n$-tuples of null-homotopic curves in the complement of $\operatorname{Fix}(\tau)$. The disks $\Delta_{i}$ lift as well, and are bounded by the lifts of the $\gamma_{i}$. Finally, since surgery on the $\gamma_{i}$ in $\overline{Y}$ gave $\overline{B^{3}}$ with fixed set consisting of the trivial arc $\overline{\ell}$, we conclude that the lifts of these surgeries equivariantly surger $Y$ to $B^{3}$. This concludes the proof of Theorem 4.2. ∎
We may also leverage the construction in this proof to obtain a concordance statement about 2-knots for whom we can construct an equivariant Seifert 3-manifold $Y$ with fixed arc $A$ such that $\operatorname{Arf}(\overline{A},\overline{Y})=1$.
@@ 4.4. Lifting the surgeries
Let $S$ be a $d$-periodic 2-knot with equivariant Seifert 3-manifold $Y$ with fixed arc $A$. If $\alpha(S,Y,A)=1$, then $S$ is equivariantly concordant to the $d$-twist spun trefoil. In particular, if $d$ is odd, then $S$ bounds an equivariant homology ball.
@@
By the construction above and the fact that every knot is band pass equivalent to either the unknot or the trefoil, we see that for any triple $S,Y,A$, with $\operatorname{Arf}(\overline{A},\overline{Y})=1$, we may ambiently surger $\overline{Y}$ to $B^{3}$ and $\overline{A}$ to the trefoil without introducing new components of the fixed set. Zeeman [Zee65] showed that the $d$-twist spin of any classical knot $K$ bounds a punctured copy of the $d$-fold cyclic cover of $S^{3}$ branched over $K$. Thus the quotient of such a Seifert solid by the rotation action permuting the twists retrieves $B^{3}$ branched along the tangle $K$. Hence, the surgeries then lift to a sequence of surgeries from $Y$ to the $d$-fold branched cover of $B^{3}$ over the trefoil, which is bounded by the $d$-twist spun trefoil.
For the latter statement, we modify this construction slightly. For integers $p,q,$ and $d$ are pairwise coprime, the Brieskorn homology sphere $\Sigma(p,q,d)$ is the $d$-fold branched cover of $S^{3}$ over the torus knot $T(p,q)$. The knot $T(p,q)$ has Arf invariant 1 if and only if $p$ is even and $q\equiv\pm 3$ mod 8; this follows from work of Levine [Lev66]. Thus if $\tau$ has order $d$ and $\operatorname{Arf}(\overline{A},\overline{Y})=1$, then we may choose an appropriate $p$ and $q$ such that $(\overline{A},\overline{Y})$ can be surgered to $(\overline{T},\overline{B^{3}})$, where $\overline{T}$ denotes the $T(p,q)$-tangle in the quotient $\overline{B^{5}}$. After this surgery, we may take the $d$-fold branched cover to see that $S$ bounds a copy of $\Sigma(p,q,d)$. ∎
At this point, we have shown that there are at most two equivariant concordance classes of $d$-periodic 2-knots: one containing the unknot, and the other containing the $d$-twist spun trefoil. However, we have not yet shown that the quantity $\alpha$ defined in Definition 4.1 is a concordance invariant of periodic 2-knots, and so we cannot say yet that these two concordance classes are not the same. In the following section, we will prove these remaining statements.
5. An obstruction to equivariant sliceness
The goal of this section is to make use of Lemma 3.5 to obstruct the existence of equivariant slice balls for 2-knots for which there exists a quotient Seifert 3-manifold with branch sets with Arf invariant 1. The first step to this is, in fact, showing the existence of $\tau$-equivariant Seifert 4-manifolds for simply periodic 3-manifolds in $S^{5}$.
5.1. Equivariant Seifert 4-manifolds
Proposition 5.1.
Let $Y$ be a simply $d$-periodic 3-manifold equivariantly embedded in $S^{5}$ with respect to the standard order $d$ rotation action $\tau$ fixing an embedded copy of $S^{3}$ such that $K:=Y\cap\operatorname{Fix}(\tau)$ is a copy of $S^{1}$. Suppose there exists an equivariant Morse function $h$ on $S^{5}$ such that $Y$ is in ordered equivariant Morse position with respect to $h$, and $h|_{K}$ is Morse and has exactly two critical points and is fully vertical with respect to $h$ away from these critical points. Then $Y$ bounds an equivariantly embedded simply $d$-periodic 4-manifold $X\subset S^{5}$ such that $X\cap\operatorname{Fix}(\tau)$ is a characteristic disk $D$ properly embedded in $X$ with $D\cdot D=0$.
Proof.
We follow roughly the same strategy as in the proof of Proposition 2.1. That is, we build $X$ handle by handle according to an equivariant handle decomposition of $Y$ coming from $h$. We do this, as before, by finding equivariant fillings for $t$-slices of $Y$, which themselves are equivariant surfaces. To better visualize these surfaces, we parametrize a neighborhood of $Y$ so that $\operatorname{Fix}(\tau)$ appears locally as the $(z,w,t)$-subspace, and so that $h$ is just projection to the $t$ coordinate. Note that the lowest minimum (maximum) of $Y$ is necessarily also the minimum (resp. maximum) of $K$, and so every $t$-slice has at least one fixed point in it. The assumption that $K$ is fully vertical with respect to $h$ away from its critical points implies that no $t$-slice has more than two fixed points in it. Thus in each $t$-slice apart from those containing the very lowest minimum and the very highest maximum, $Y$ appears as a surface in 4-space, equivariant with respect to rotation about the $(z,w)$-axis, with two fixed points. Our goal is therefore to construct smoothly varying 3-dimensional fillings $X_{t}$ of these surfaces $Y_{t}$ such that each of these slices intersects $\operatorname{Fix}(\tau)$ in a single smoothly varying arc. These arcs stacked on top of each other will then form a disk, which will be $\operatorname{Fix}(\tau)\cap X$ as desired. As before, we provide a schematic showing the handle attachments at different stages.
[…]
Figure 10. A schematic showing the attachment of handles of $X$ with respect to the Morse function $h$.
$\boldsymbol{f=0.}$ Since $Y$ is in equivariant Morse position, we can assume that all of its 0-handles occur in $d$-tuples, except for one, which is equivariant with respect to $\tau$. We may begin constructing $X$ by filling in these 0-handles with 4-dimensional index 0 right half-handles. This completes $X$ up to $t=\frac{1}{2}$.
$\boldsymbol{f=\frac{1}{2}.}$ At this point, $X_{\nicefrac{{1}}{{2}}}$ is an equivariant collection of 3-balls in ambient 4-space. As in the previous case, we want to fill the 1-handles of $Y$ with 4-dimensional right half-handles. However, the trajectories from the index 1 critical points of $Y$ may link slices of the 0-handles of $Y$. As before, we consider these trajectories intersecting the 0-handles of $Y$ in $h^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$. To avoid these intersections, we preemptively attach interior 1-handles to $X$ at $t=\frac{1}{2}-\epsilon$ along the trajectories to create 3-dimensional tubes for these trajectories to pass through.
$\boldsymbol{f=1.}$ After these interior 1-handles are attached, we may proceed to attach the 3-dimensional 1-handles and their 4-dimensional fillings.
$\boldsymbol{f=\frac{3}{2}.}$ Following the same general strategy, we now examine the descending manifolds of the 3-dimensional index 2 critical points at $t=2$. Because $Y\cap\operatorname{Fix}(\tau)$ is precisely $K$, we know none of these descending manifolds intersect $\operatorname{Fix}(\tau)$, as all such intersections are already accounted for by our assumption that $K$ has precisely two critical points with respect to the ambient Morse function. Thus we may safely attach right half-handles to $X$ for each 2-handle of $Y$ without fear of breaking equivariance or introducing new components of the fixed set. However, the index 2 descending manifolds may intersect interior 1-handles of $X$ that we have already attached. As before, suppose that all of these intersections occur within the preimage $h^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
To avoid these intersections, note that the generic intersection between a 2-handle core and a 4-dimensional 1-handle in $B^{5}$ is a disk in the 2-handle core. Thus we can use this disk as a guide to preemptively attach 4-dimensional interior 2-handles to compress the 4-dimensional 1-handles before attaching the 3-dimensional 2-handle. This is completely analogous to the preemptive 1-handles we attached to make way for the 1-handles of $Y$. Note further that because the disks we use as guides for attaching our 4-dimensional 2-handles are already embedded in $B^{5}$ as the cores of existing 3-dimensional 2-handles, they have trivial normal bundles. Thus from the perspective of $X$, the compressions are achieved by attaching interior 2-handles, with framings given by a choice of trivial disk sub-bundle of the normal bundle. Furthermore, since none of the 2-handles of $Y$ intersect $\operatorname{Fix}(\tau)$, we may arrange so that the parallel 2-handles of $X$ do not intersect $\operatorname{Fix}(\tau)$ either.
$\boldsymbol{f=2.}$ Once we have performed these compressions at $t=\frac{3}{2}-\epsilon$, we may safely proceed upwards in $t$ and attach the 3-dimensional 2-handles of $Y$ and their 4-dimensional half-handle fillings at $t=2$.
$\boldsymbol{f=\frac{5}{2}.}$ At this point, for $t\in(2+\epsilon,\frac{5}{2}-\epsilon)$, $Y_{t}$ appears as an equivariant unlink of unknotted spheres with a pair of fixed points on a single component, and $X_{t}$ appears as a connected equivariant 3-manifold with boundary $Y_{t}$ and fixed set a single, “straight” line (with respect to our parameterizing coordinates). Now every component of $Y_{\nicefrac{{5}}{{2}}-\epsilon}$ bounds a 3-ball embedded in $B^{5}_{\geq 2+\epsilon}$ given by the descending manifolds of each maximum of $h$ on $Y$. As in the 3-manifold case, we may isotope these 3-ball caps straight down into the $t=\frac{5}{2}-\epsilon$-slice, using their boundary unknotted 2-spheres as guides near the boundary. We may also assume that these 3-balls are all equivariant, meaning that one central 3-ball intersects $\operatorname{Fix}(\tau)$ in a straight arc connecting the fixed points of $Y_{2+\epsilon}$ and the rest are permuted by $\tau$.
As in the lower-dimensional case, the boundaries of tubular neighborhoods of each of these 3-balls then form splitting spheres for $Y_{\nicefrac{{5}}{{2}}-\epsilon}$. These splitting spheres intersect $X_{\nicefrac{{5}}{{2}}-\epsilon}$ in a collection of 3-balls, which we may then use as equivariant compressing 3-balls for $X_{\nicefrac{{5}}{{2}}-\epsilon}$. We can then equivariantly compress $X$ along these 3-balls by attaching interior 3-handles to $X_{\nicefrac{{5}}{{2}}}$ and filling in closed components.
$\boldsymbol{f=3.}$ In fact, because the splitting spheres we found were the boundaries of small neighborhoods of balls bounded by components of $Y_{\nicefrac{{5}}{{2}}}$, each component of $X_{\nicefrac{{5}}{{2}}}$ with boundary is itself a 3-ball. We may subsequently fill the remaining closed components with equivariant 4-manifold fillings. Finally, we trace the isotopy of the 3-ball caps up to the maxima of $Y$ to find equivariant 4-balls that fill these maxima. This completes the construction of $X$ and the proof of Proposition 5.1. ∎
Remark 5.2.
The approach used in Propositions 2.1 and 5.1 is indicative of a more general strategy for building embedded Seifert $n$-manifolds inside $B^{n+1}$ explicitly handle by handle. That is, for every $k$-handle attachment of the boundary, attach interior $k$-handles to the Seifert filling to make way for the descending manifolds of the corresponding critical points. This approach was necessary for this paper in order to control the intersection of the the Seifert manifolds with the ambient fixed set.
5.2. Proof of obstruction
Recall that for a periodic 2-knot $S$ with equivariant Seifert 3-manifold $Y$ such that $Y\cap\operatorname{Fix}(\tau)$ is an arc $A$, the quantity $\alpha(S,Y,A)$ was defined to be:
$\alpha(S,Y,A)=\operatorname{Arf}(\overline{A},\overline{Y})$
where $\overline{A}$ and $\overline{Y}$ denote the image of these manifolds after taking the quotient by $\tau$.
Theorem 5.3.
Let $S$ be a periodic 2-knot and suppose $Y_{1}$ and $Y_{2}$ are two quotients of equivariant Seifert 3-manifolds with fixed arcs $A_{1}$ and $A_{2}$, respectively. Then $\alpha(S,Y_{1},A_{1})=\alpha(S,Y_{2},A_{2})$.
Proof.
Begin by embedding $S$ in $S^{4}$, thought of as an equator of $S^{5}$, so that the $S^{4}$ bounds two equivariant 5-balls whose union is all of $S^{5}$. Note that $\tau$ can then be extended to $S^{5}$ by gluing the actions on each $B^{5}$ on either side of the equator together, so that the action on $S^{5}$ fixes a 3-sphere. Parametrize a neighborhood of $S$ in $S^{5}$ with our usual conventions, though we now allow for $t$ to take negative values in order to describe points in both 5-balls.
Embed $Y_{1}$ in the $t\geq 0$ copy of $B^{5}$ such that $A_{1}$ has exactly one maximum with respect to $t$. In the other copy of $B^{5}$, embed $Y_{2}$ such that $A_{2}$ has exactly one minimum with respect to $t$. Note that these embeddings can be found as a result of Lemma 4.3.
The union $Y$ of $Y_{1}$ and $Y_{2}$ along $S$ is then a simply periodic closed 3-manifold, and its fixed set under $\tau$ is the union $K$ of $A_{1}$ with $A_{2}$. Note that with respect to projection onto the $t$-axis, $K$ has exactly two critical points. The pair $Y,K$ then satisfies the conditions of Theorem 5.1. Therefore, $Y$ bounds an equivariant, spin Seifert 4-manifold $X$ with fixed set a characteristic, 0-framed disk $D$ with boundary $K$.
We now take the quotient of this whole construction by $\tau$. Note that $\overline{D}$ remains characteristic and 0-framed in $\overline{X}$. Thus by Corollary 3.7, $\operatorname{Arf}(\overline{K},\overline{Y})$ must vanish. Hence by the additivity of the Arf invariant proved in Lemma 3.9, we conclude that $\alpha(S,Y_{1},A_{1})=\alpha(S,Y_{2},A_{2})$. ∎
Having concluded this, we see that $\alpha$ is independent of the choice of equivariant Seifert solid $Y$, and so it is an invariant of periodic 2-knots. Thus by Theorem 4.2 and Proposition 4.6, we have shown that the invariant $\alpha$ completely classifies $d$-periodic 2-knots up to equivariant concordance. As an immediate corollary, we obtain the following.
Theorem 5.4.
For all $d\geq 2$, $\mathcal{C}^{2}_{d}\cong\mathbb{Z}/2\mathbb{Z}$, generated by the $d$-twist spun trefoil.
Proof.
Consider the map $\varphi$ from $\mathcal{C}^{2}_{d}$ to $\mathbb{Z}/2\mathbb{Z}$ given by $S]\mapsto\alpha(S)$. By Theorem 5.3, $\alpha$ is well-defined on equivariant concordance classes, and so $\varphi$ is a well-defined map. The combination of Theorem 4.2 and Proposition 4.6 show that $\varphi$ is injective, as we can realize an equivariant concordance between any $d$-periodic 2-knot and either the unknot or the $d$-twist spun trefoil. We have also seen in the proof of Proposition 4.6 that $\alpha$, and thus $\varphi$, takes the value 1 on the $d$-twist spun trefoil for all $d$.
Finally, to show that $\varphi$ is indeed a homomorphism, we appeal to the additivity of the Arf invariant. Let $S$ and $S^{\prime}$ be two periodic 2-knots with equivariant Seifert 3-manifolds $Y$ and $Y^{\prime}$. Then their equivariant connected sum $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ bounds an equivariant Seifert 3-manifold $Y\natural_{\mathbb{Z}/d\mathbb{Z}}Y$ whose fixed set is the connected sum of the fixed sets of $Y$ and $Y^{\prime}$ as arcs. By Lemma 3.9, the Arf invariant of the connected sum of the fixed sets is the sum of the individual Arf invariants mod 2. Thus $\alpha(S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime})=\alpha(S)+\alpha(S^{\prime})$. This proves that $\varphi$ is an isomorphism, and thus that $\mathcal{C}^{2}_{d}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. ∎

### v2

@@ 6.2. The equivariant concordance invariant
Let $\tilde{Y}_{1}$ and $\tilde{Y}_{2}$ be two equivariant Seifert manifolds for $\tilde{S}$, and suppose that each intersects $\operatorname{Fix}(\rho)$ in arcs $\hat{A}_{1}$ and $\hat{A}_{2}$, respectively. Let $Y_{1}$ and $Y_{2}$ be their quotients by $\rho$, intersecting $\operatorname{Branch}(\rho)$ in arcs $A_{1}$ and $A_{2}$, respectively. As in the proof of Proposition 5.4, parametrize a neighborhood $U$ of $S$ in $S^{5}$ with coordinates $(x,y,z,w,t)$ such that $\operatorname{Branch}(\rho)\cap U$ consists of the $(z,w,t)$-subspace, and that $S^{4}\cap U$ is the $t=0$ subspace. We can then apply Lemma 5.3 to isotope $Y_{1}$ and $Y_{2}$ rel. boundary into opposite copies $B_{1}$ and $B_{2}$ of $B^{5}$ bounded by $S^{4}$ such that each intersects $\operatorname{Branch}(\rho)$ in an arc, which we continue to denote $A_{1}$ and $A_{2}$. The lemma also provides us with a Morse function $f$, which we can take to be projection onto the $t$ coordinate, such that each $A_{i}$ is fully vertical with respect to $f|_{Y_{i}}$ away from a single critical point. By Proposition 3.5, we can isotope each $Y_{i}$ rel. boundary to ordered Morse position with respect to $f$ through isotopies supported away from the $A_{i}$’s, ensuring that they will lift to isotopies in the branched cover as in the proof of Proposition 3.11.
Let $Y$ be the union of $Y_{1}$ and $Y_{2}$ along $S$, and let $K$ be the union of $A_{1}$ and $A_{2}$ along their boundary. Then $Y,A,$ and $\operatorname{Branch}(\rho)$ satisfy the conditions of Proposition 4.3, and so we have that $\operatorname{Arf}(K,Y,\mathfrak{s}_{0})=0$ for $\mathfrak{s}_{0}$ the unique spin structure on $S^{5}$. Conclude by the additivity of the Arf invariant in Lemma 2.14 that $\operatorname{Arf}(A_{1},Y_{1},\mathfrak{s}_{0})=\operatorname{Arf}(A_{2},Y_{2},\mathfrak{s}_{0})$. ∎
As in Proposition 5.5, we use the same trick to show that $\alpha_{d}$ vanishes on $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice periodic 2-knots.
Proposition 6.6.
Let $\tilde{S}$ be a $d$-periodic 2-knot in $\tilde{S}^{4}$ and suppose that $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice. Then $\alpha_{d}(\tilde{S})=0$.
Proof.
Parametrize a neighborhood of the quotient sphere $S$ in $S^{5}$ with coordinates $(x,y,z,w,t)$ such that $S$ lies in the $t=0$ hypersurface, and such that $\operatorname{Branch}(\rho)$ consists of the $(z,w,t)$-subspace. Let $f$ be projection onto the $t$-axis.
Since $\tilde{S}$ is equivariantly slice, there exists an equivariant concordance $\tilde{C}\subset\tilde{S}^{4}\times I$ cobounded by $\tilde{S}$ and the unknot $\tilde{U}$. Recall that our definition of equivariantly slice requires this embedding to be level-preserving on the fixed sets of these actions. We may therefore embed the quotient concordance inside $S^{4}\times I$ into our parametrized neighborhood of $S^{5}$ in such a way that $C\cap\operatorname{Fix}(\rho)$ is fully vertical with respect to $f|_{C}$. We may then cap off this quotient concordance with a spanning 3-ball $B$ bounded by the quotient unknot $U$ in the $t=1$ slice, pushed slightly deeper in the direction of increasing $t$ using Lemma 5.3 such that its intersection with $\operatorname{Branch}(\rho)$ is fully vertical with respect to $f|_{B}$, to obtain a slice ball $B^{\prime}$ for $S$ which intersects $\operatorname{Branch}(\rho)$ in an unknotted arc $\ell$ which is fully vertical with respect to $f|_{B^{\prime}}$.
Now let $\tilde{Y}$ be an equivariant Seifert 3-manifold for $\tilde{S}$ intersecting $\operatorname{Branch}(\rho)$ in a null-homologous arc, and let $Y$ be its quotient under $\rho$. Again by Lemma 5.3, we may isotope $Y$ into the $t\leq 0$ 5-ball such that its intersection with $\operatorname{Branch}(\rho)$ in an arc $A$ which is fully vertical with respect to $f|_{Y}$. Then the closed 3-manifold $Y\cup B^{\prime}$ together with $A\cup\ell$ satisfies the conditions of Proposition 4.3, and so we have that $\operatorname{Arf}(A\cup\ell,Y\cup B^{\prime},\mathfrak{s}_{0})=0$. Again by the additivity established in Lemma 2.14, this implies $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=\operatorname{Arf}(\ell,B^{\prime},\mathfrak{s}_{0})$, which then vanishes because $\ell$ is unknotted and $B^{\prime}$ is a 3-ball. ∎
We can also establish additivity for $\alpha_{d}$.
Lemma 6.7.
Let $\tilde{S}$ and $\tilde{S}^{\prime}$ be two $d$-periodic 2-knots inside two copies of the 4-sphere which equipped with $\mathbb{Z}/d\mathbb{Z}$ actions $\rho$ and $\rho^{\prime}$, respectively. Then
@@
Proof.
Let $S$ and $S^{\prime}$ be the quotient spheres of $\tilde{S}$ and $\tilde{S}^{\prime}$. Let $Y$ and $Y^{\prime}$ be Seifert 3-manifolds for the corresponding spheres $S$ and $S^{\prime}$ and let $A$ and $A^{\prime}$ be null-homologous arcs of intersection of $Y,Y^{\prime}$ with $\operatorname{Branch}(\rho)$ and $\operatorname{Branch}(\rho^{\prime})$, respectively. Then by the equivariant connected sum given in Definition 1.12, we see that the partial boundary connected sum $Y$ of $Y$ and $Y^{\prime}$ forms a Seifert 3-manifold for the quotient sphere of $\tilde{S}\#_{\mathbb{Z}/d\mathbb{Z}}\tilde{S}^{\prime}$ which intersects the connected sum $\operatorname{Branch}(\rho)\#\operatorname{Branch}(\rho^{\prime})$ in the boundary connected sum of the tangles $A$ and $A^{\prime}$. The result then follows from Lemma 2.14. ∎
6.3. Building equivariant concordances
@@
Let $\tilde{S}$ be a $d$-periodic knot with $\alpha_{d}(\tilde{S})=0$. Then $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice.
@@
Let $\tilde{Y}$ be an equivariant Seifert 3-manifold for $\tilde{S}$ intersecting $\operatorname{Fix}(\rho)$ in a null-homologous arc $\hat{A}$. Let $S$ be the quotient sphere for $\tilde{S}$, and note that $\tilde{Y}$ descends to a Seifert 3-manifold $Y$ for $S$ intersecting $\operatorname{Branch}(\rho)$ in a null-homologous arc $A$. By Lemma 5.3, we may isotope $Y$ rel. boundary into the interior of $B^{5}$ while keeping its intersection with $\operatorname{Branch}(\rho)$ still a single arc, which we continue to call $A$. We are now in precisely the same setup as in Subsection 5.2, taking $\operatorname{Branch}(\rho)$ to be the removed sphere. We may therefore apply Lemma 5.10 to find a sequence of ambient surgeries taking $Y$ to a 3-ball $B$ bounded by $S$ which do not introduce any new intersections with $\operatorname{Branch}(\rho)$.
As in Proposition 5.12, the condition that $\alpha_{d}(S)=0$ requires that $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=0$, which by Proposition 5.9 means that $\operatorname{Arf}(A,B,\mathfrak{s}_{0})=0$. We can then apply Lemma 5.11 to find another sequence of ambient surgeries on $B$ turning it into another 3-ball $B^{\prime}$ such that $B^{\prime}\cap\operatorname{Branch}(\rho)$ is an unknotted arc $\ell$. Since the disks surgered along did not intersect the branch set of $\rho$, we may lift the surgery curves to a set of equivariant curves on $\tilde{Y}$. The framed disks bounded by these curves lift as well, and so we obtain from this procedure a set of equivariant ambient surgeries converting $\tilde{Y}$ into an equivariant 3-ball bounded by $\tilde{S}$. ∎
Proof of Theorem 6.1.
The proof is exactly the same as in Theorem 5.2. Propositions 6.6 and 5.12 show that a $d$-periodic 2-knot $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice if and only if $\alpha_{d}(\tilde{S})=0.$ Since $\alpha_{d}$ is additive under equivariant connected sum by Proposition 6.7 and valued in $\mathbb{Z}/2\mathbb{Z}$, every periodic 2-knot which is not $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice is the inverse of every other such 2-knot in the corresponding equivariant concordance group $\mathcal{C}_{2}^{d}$. Hence all such 2-knots must be equal and have order 2 in $\mathcal{C}_{2}^{d}$, which thus must be isomorphic to $\mathbb{Z}/2\mathbb{Z}$. ∎